OpenStudy (anonymous):

29.2% HCL stock solution has a density 1.25g/ml the MM of HCL is 36.5 . THe volume (Ml) of the stock solution required to prepare 200Ml of solution of 0.4M HCl is?

OpenStudy (anonymous):

29.2% is mass percentage?

OpenStudy (anonymous):

Yes (W/W)

OpenStudy (anonymous):

0.4M = 0.4mol/dm^3 200 ml = 0.2 dm^3 moles of HCL in required solution is \[0.4 \times 0.2 \text{ }mol \text{ } = 0.08 \text{ }mol\] convert to mass as mol = mass/Mr \[mass = 36.5 \times 0.08 =2.92g\] now we just need to work out how much stock solution is 2.92g

OpenStudy (anonymous):

2.92g = 29.2% \[2.92g =0.292x\] where x is the mass of the whole stock solution we need to add. so \[x = \frac{2.92}{0.292} = 10g\] now we need to use the density of the stock solution to work out the olume containing 10g does this make sense to you so far?

OpenStudy (anonymous):

volume*

OpenStudy (anonymous):

ok @eigenschmeigen plz go on

OpenStudy (anonymous):

@eigenschmeigen the answer should be 8

OpenStudy (anonymous):

we need 10g with 1.25g/ml density = mass/volume volume = mass/density volume = 10/1.25 = 8ml